The cooling system of an automobile engine holds 16 liters of fluid. The system is filled with a mixture of 80% water and 20% antifreeze. It is necessary to increase the amount of antifreeze to 40%. This is to be done by draining some of the mixture and adding pure antifreeze to bring the total amount of fluid back up to 16 liters. How much of the mixture should be drained and replaced with pure antifreeze?
Answer by TimothyLamb(4379)

x = liters of pure (100%) antifreeze to add, after draining x liters of original coolant
---
0.20*(16 - x) + 1.00x = 0.40*16
0.20*16 - 0.20x + x = 0.40*16
x - 0.20x = 0.40*16 - 0.20*16
0.80x = 0.40*16 - 0.20*16
x = (0.40*16 - 0.20*16)/0.80
(0.40*16 - 0.20*16)/0.80
---
copy and paste the above expression in to this calculator:
http://sooeet.com/math/scientific-calculator.php
---
x = 4
---
Free algebra tutoring live chat:
http://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
http://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
http://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
http://sooeet.com/math/system-of-linear-equations-solver.php
---